OpenStudy (anonymous):

How do you solve this? Consider the functions f(x) = 2(x − 3) and g(x) = 3x. What is the value of (f + g)(3)?

hero (hero):

(f + g)(x) = f(x) + g(x)

hero (hero):

f(x) + g(x) = 2(x - 3) + 3x = 2x - 6 + 3x = 2x + 3x - 6 = 5x - 6 f(3) + g(3) = 5(3) - 6

OpenStudy (anonymous):

The 3 do I just plug it in to both equations then add the two equations solutions?

hero (hero):

You could do it that way. Another way is what I have shown above.

hero (hero):

You should learn both methods.

OpenStudy (anonymous):

Alright well, may I ask you a similar question in this "question" or would you like me to create a whole new thing? It is on the same topic.